The sum of two numbers is 38. Their difference is 10. What are the two numbers?

Answer by gems(12)

x+y=38
x-y=10
add the two equations
2x=48
x=24
if x= 24, subst. to x+y=38
24+y=38
y=38-24
y=14
therefore x=24, y=14